Grade 99 grade maths

if a and b are the roots of the qudratic equation 5y2-7y+1=0 the find the vlue of 1/a+1/b

himanshu
8 Years agoGrade 9

1 Answer

Prashant
8 Years ago
Simply sove 5y^2+7y+1, u`ll get a=7/10+√29/10 and b=7/10-√29/10Then, simply reciprocal a and b.U`ll get 1/a+1/b=140/20=7.I hope u`ll get it.